# What are they talking about? Benchmarking Large Language Models for Knowledge-Grounded Discussion Summarization

### v1

@@ 4.4 Annotation and Evaluation for ABS
Evaluates ABS overall performance by comprehensively considering both coverage and focus:
4.5 Annotation and Evaluation for AOS
Clear Atomic Opinions Annotation.
The annotations are created entirely by experts and involve three steps: main opinion extraction, implicit reference clarification, and clear opinion atomization. First, we require experts to extract the main opinions from their respective utterances in the discussion. Next, we ask them to identify referential pronouns and phrases within these opinions and clarify them (i.e., through anaphora resolution or information supplementation) using referents (e.g., entities, facts, or events) from the news article to produce clear opinions. Finally, we instruct the experts to decompose these clear opinions into atomic opinion units, adhering to the principles of indivisibility and independence (see Appendix B for detailed annotation statistics).
Fine-Grained Quantification Metric.
Let be the LLM-generated AOS, and be the set of clear atomic opinions. The LLM-based opinion verification function is defined as:
where and indicate whether the opinion can be inferred from .
Metric . CAO Recall
Assesses AOS coverage of clear atomic opinions:
Fine-Grained Error Detection.
Unlike query-focused ABS, the integrativity of AOS determines that it is almost impossible to exhaust the space of erroneous atomic opinions. As a result, defining precision and F1-score for AOS evaluation is impractical. However, we can identify why certain opinions are not covered by AOS through fine-grained error detection. Specifically, for opinions that cannot be inferred from AOS, we define five error types and conduct LLM-based error classification: implicit reference unclarified, implicit reference incorrectly clarified, opinion misattribution, opinion fact inconsistency, and opinion sentiment distortion (detailed error definitions are provided in Appendix D.7).
4.6 Overall Performance Evaluation
We use the Geometric Mean to assess pattern-level overall performance for two reasons: (1) Its multiplicative property reflects the joint fulfillment rate of the background and opinion summaries, emphasizing the holistic synergy between these two components. (2) The root normalization ensures the dimensional consistency of the component metrics. Specifically, we define the overall performance metrics for the EBS-AOS and ABS-AOS patterns as:
@@ 4.6 Overall Performance Evaluation
| EBS-AOS Pattern | ABS-AOS Pattern | |||||
| BSP (R-P-F1) | CAO (R) | OP (GM) | KBSAF (R-P-F1) | CAO (R) | OP (GM) | |
| GPT-4o | 73.39↑1.56-88.10↑0.34-78.12↑1.27 | 76.18↑0.54 | 76.24↑1.04 | 63.57↓1.08-61.73↑1.62-58.34↑0.60 | 69.42↓0.21 | 61.09↑0.47 |
| GPT-4-turbo | 73.03↓7.72-87.42↑1.67-77.14↓3.80 | 72.73↑0.45 | 73.94↓1.47 | 46.28↓5.78-54.42↑5.33-47.26↓1.64 | 58.32↓1.90 | 48.28↓0.74 |
| GPT-4o-mini | 76.23↓0.30-67.71↑0.31-67.92↓0.03 | 29.08↓1.00 | 34.24↓0.97 | 33.75↑0.48-45.07↑0.52-36.51↑0.52 | 27.74↓0.41 | 24.64↓0.21 |
| Claude 3 Opus | 65.60↑2.21-85.01↓4.57-72.07↓1.31 | 75.20↓2.19 | 72.09↓2.45 | 51.10↑1.03-74.59↓1.13-58.03↓0.91 | 69.95↓2.41 | 60.72↓1.29 |
| Claude 3.5 Sonnet | 80.36↑2.42-87.93↓0.62-82.33↑0.46 | 74.93↓5.30 | 77.12↓4.45 | 40.74↑5.84-65.18↓0.79-47.75↑3.51 | 58.55↑1.29 | 48.83↑2.60 |
| Claude 3.5 Haiku | 62.69↓8.39-76.01↑0.37-66.02↓5.66 | 40.35↓9.13 | 46.02↓10.1 | 37.42↓6.90-48.72↓4.65-39.91↓6.63 | 29.64↓5.71 | 27.22↓4.48 |
| Gemini 1.5 Pro | 84.64↓4.38-82.25↑4.43-79.73↑1.36 | 76.86↓0.53 | 76.71↑0.49 | 50.40↓4.13-52.33↑0.28-48.42↓2.11 | 69.09↓6.08 | 54.83↓4.45 |
| Llama-3.1-405B | 79.09↑2.72-77.96↓1.60-75.16↑0.36 | 64.25↓0.01 | 67.58↑0.01 | 38.19↓1.08-58.47↑7.38-43.39↑1.12 | 52.98↓2.44 | 43.79↑0.11 |
| Mistral Large | 68.81↑1.40-78.55↓0.18-71.07↑0.89 | 60.82↓2.62 | 63.63↓1.59 | 53.91↓3.05-56.78↓0.32-52.57↓1.94 | 46.24↓1.71 | 46.36↓2.26 |
| DeepSeek-V3 | 86.98↓1.19-73.83↑1.37-75.66↑1.07 | 64.98↑1.84 | 66.50↑1.61 | 47.64↑0.65-42.12↑0.60-42.17↑0.65 | 56.02↑0.77 | 44.09↑0.72 |
| Qwen-Max | 74.59↓5.53-79.86↓1.30-73.79↓2.78 | 60.93↓1.24 | 64.34↓1.59 | 46.49↓0.02-53.59↓0.35-46.87↓0.17 | 45.29↓1.54 | 40.87↓0.23 |
| GLM-4-Plus | 80.28↓1.03-72.81↑0.61-71.17↑0.34 | 69.34↓1.73 | 67.55↓0.43 | 43.50↓1.77-46.93↑0.43-41.64↓0.61 | 55.80↓3.17 | 43.54↓1.26 |
5 Evaluation Setup
Evaluated LLMs.
We evaluate leading LLMs, covering both the most advanced and lightweight variants555Our evaluation began on January, 2025, and all LLMs used the latest official API versions available at that time.: GPT-4o, GPT-4-turbo, GPT-4o-mini, Claude 3 Opus, Claude 3.5 Sonnet, Claude 3.5 Haiku, Gemini 1.5 Pro, Llama-3.1-405B, Mistral Large, DeepSeek-V3, Qwen-Max, GLM-4-Plus. All model sources are listed in Appendix E.
Solution Paradigms.
We benchmark LLMs for KGDS under two interaction paradigms: (1) Single-turn Structured-prompt: A standard prompt with well-structured Li et al. (2023) contains input content, input definition, task description, output definition, and return format (in Appendix D.3). (2) Multi-turn Self-reflection: A second-round self-reflection Shinn et al. (2023) instruction with step-by-step chain-of-thought Wei et al. (2022) reasoning after the structured prompt (in Appendix D.4).
@@ 6 Analysis
In this part, we primarily analyze the commonalities and differences among LLMs in KGDS. Sections §6.1, §6.2, and §6.3 respectively discuss the performance of background summary, opinion summary, and both patterns under structured-prompt. §6.4 explores the impact of LLM self-reflection.
6.1 Background Summary Analysis
@@
From Table 1, we find that most LLMs achieve moderate retrieval performance () and lightweight models (i.e., GPT-4o-mini and Claude 3.5 Haiku) perform poorly. However, from and , we observe significant polarization among LLMs, which indicates distinct retrieval strategies: some prioritize precision at the cost of recall (e.g., GPT-4-turbo), while others do the exact opposite (e.g., DeepSeek-V3). Such imbalance reveals that the current LLMs struggle with the precision-recall trade-off. Moreover, we investigate the paragraph retrieval ratio (Figure 2) and identify that most LLMs exhibit either under- or over-retrieval, which is consistent with imbalance.
LLMs prefer expert consistently annotated EBS over their own extractions.
We conduct an investigation to verify whether our EBS is the true gold standard. Specifically, for EBS extracted by LLM that is inconsistent with expert-annotated, we instruct the LLM to perform a second-round feedback judgment to determine which EBS is more aligned with its definition (the instruction is provided in Appendix D.5). As shown in Figure 3, all LLMs unanimously and overwhelmingly prefer our annotations, with the average win rate of , indicating that: (1) Our expert-annotated EBS can be regarded as reliable ground truth. (2) The current LLMs possess latent awareness of optimal retrieval boundaries but struggle to implement them.
@@
As shown in Table 1, all LLMs demonstrate unsatisfactory performance (). The weak () reveals that LLMs often omit key facts, while the mediocre () indicates persistent inclusion of irrelevant facts. This dual failure reflects the fundamental deficiencies of LLMs in meeting the requirements of coverage (i.e., completeness) and focus (i.e., conciseness). Moreover, we also observe polarization among LLMs: some prioritize precision at the cost of recall (e.g., Claude 3 Opus), while others attempt to balance both (e.g., GPT-4o). Unlike EBS, we did not find any extreme recall-oriented models, indicating that LLMs tend to be either conservative or balanced in ABS generation.
LLMs are better retrievers than generators for background summary.
As presented in Table 1, all LLMs exhibit higher BSP metrics compared to their KBSAF counterparts777BSP and KBSAF metrics are directly comparable, as paragraph and atomic fact differ only in granularity, and there is no difference in mathematical form between the twin metrics.. This gap between retrieval and generation indicates that LLMs possess stronger in-context recognition capabilities for coarse-grained paragraphs than fine-grained facts.
6.2 Opinion Summary Analysis
@@
From Table 1, we find that decreases as the background summary quality declines (i.e., ) across all LLMs during pattern transitions. Additionally, as shown in Figures 4 and 5, is highly positively correlated with and among models in independent patterns. These findings indicate that a high-quality background summary facilitates knowledge integration in AOS generation. Meanwhile, individual differences suggest that the model-inherent integration ability is also a key factor affecting AOS quality. For instance, Gemini 1.5 Pro and Claude 3.5 Sonnet show relatively stronger (i.e., above the line in Figure 5) and weaker (i.e., below the line in Figure 4) integration capabilities in ABS-AOS and EBS-AOS patterns, respectively.
As for absolute performance, the most advanced models (highlighted LLMs in Table 1) only achieve an average of across both patterns, indicating that numerous opinions are being incorrectly integrated. Moreover, we observe that lightweight models (i.e., GPT-4o-mini and Claude 3.5 Haiku) exhibit significant flaws in , suggesting that AOS integration is sensitive to parameter scale.
AOS error detection analysis.
As shown in Table 2, a significant proportion of integration errors are concentrated in IRU and IRIC across all LLMs in both patterns, indicating that LLMs struggle to effectively and correctly clarify implicit references during AOS integration. Furthermore, although the ratios of OFI, OSD, and OM errors are relatively lower, their presence still emphasizes the inherent deficiencies of factual inconsistency, sentiment distortion, and incorrect speaker assignment in LLMs.
@@
6.3 Overall Performance Analysis
Performance stratification among LLMs.
We observe that metrics exhibit distinct hierarchies in independent patterns and cross-pattern (refer to Figures 6, 8, and 9 in Appendix A). This suggests that the performance of LLMs improves in a stepwise manner rather than continuously as their intelligence advances in KGDS. Moreover, even the best-performing models achieve less than average overall performance across both patterns (see Figure 9), highlighting that KGDS remains a challenging task for current SOTA LLMs.
Cross-pattern stability of LLMs.
By analyzing the performance gap between the two patterns (see Figures 6, 7, and 8 in Appendix A), we find that different LLMs excel in different patterns. For example, Claude 3.5 Sonnet exhibits a significant gap ( ), while Claude 3 Opus demonstrates a relatively smaller gap ( ), indicating stronger cross-pattern stability.
LLMs perform better with EBS-AOS than ABS-AOS for KGDS.
As presented in Table 1, all LLMs achieve superior in EBS-AOS pattern than ABS-AOS, which is attributed to the more complete and accurate background summary and higher-quality opinion summary. Hence, we suggest prioritizing EBS-AOS in real-world KGDS for better implementation.
@@ 6.4 Self-Reflection Impact
From Table 1, we observe that the performance fluctuations (i.e., or ) after self-reflection are limited (the maximum average fluctuation across all metrics is for ). This means that self-reflection does not fundamentally influence the capacities of LLMs in KGDS, revealing two key limitations: (1) LLMs lack sufficient self-evaluation capacities for KGDS. (2) The reflection strategies of LLMs struggle to provide excellent reasoning paths for KGDS.
@@
As presented in Table 2, most LLMs show a decrease in IRIC and an increase in IRU (e.g., GPT-4o with IRIC , IRU ). This indicates that self-reflection drives LLMs to adopt more cautious strategies when clarifying implicit references — preferring to leave references unclarified rather than risk incorrect clarification. However, excessive conservatism in clarifying implicit references harms AOS quality, as evidenced by the widespread decrease in in Table 1.

### v2

@@ 4.1 Dimensions
Overall Quality: A holistic assessment of the summary, considering both coverage and focus.
Opinion Summary.
We evaluate the output performance according to the following dimensions666Unlike background summaries, opinion summaries are inherently integrative. The space of opinions that could be wrongly integrated is effectively unbounded, making it infeasible to construct an exhaustive set of all incorrect atomic opinions to be excluded. Therefore, measuring a Focus dimension is impractical. Instead, we identify integration errors in the summary for uncovered clear atomic opinions.:
Overall Quality: Whether the summary successfully integrates clarified opinions, enabling it to cover all clear atomic opinions.
Integration Error: For uncovered clear atomic opinions, identify the errors in the summary.
Summarization Paradigm.
We assess the overall quality of each summarization paradigm:
Overall Quality: How well the background and opinion summaries work together.
4.2 Methods and Metrics
EBS Evaluation. We evaluate its quality by matching the paragraph indices from the system output against our annotations of supporting and nonsupporting paragraphs. We utilize Supporting Paragraph Recall, Precision, and F1-score to measure the coverage, focus, and overall quality of the summary, respectively.
ABS Evaluation. We employ an LLM-based verifier Wei et al. (2024) to check if the summary entails our annotated atomic facts. This process determines whether the summary includes key supporting facts while excluding nonsupporting ones. We use three fine-grained metrics, Key Supporting Atomic Fact Recall, Precision, and F1-score, to quantify the three dimensions of the summary. We provide detailed formulas in Appendix C.1.
AOS Evaluation. Similar to ABS, we utilize an opinion verifier to measure its coverage of our annotated atomic opinions. We utilize Clear Atomic Opinion Recall to assess its overall quality. For each uncovered atomic opinion, we perform LLM-based error detection to identify the specific integration failure. Specifically, we define five fine-grained error types: implicit reference unclarified, implicit reference incorrectly clarified, opinion misattribution, opinion fact inconsistency, and opinion sentiment distortion. Detailed formulas and error definitions are provided in Appendix C.2 and F.6.
Paradigm-Level Quality Evaluation. A successful KGDS output requires high quality in both its background and opinion summaries. To capture this dependency, we evaluate the paradigm-level quality using the Geometric Mean of the two sub-summary overall scores. The multiplicative nature of this metric ensures a high score is only achieved when both components are strong, reflecting their joint fulfillment of the task goal. Moreover, root normalization maintains dimensional consistency between the paradigm-level score and its parts.
@@ 4.2 Methods and Metrics
| EBS-AOS Paradigm | ABS-AOS Paradigm | |||||
| SP (R-P-F1) | CAO (R) | PLQ (GM) | KSAF (R-P-F1) | CAO (R) | PLQ (GM) | |
| GPT-4o | 73.39↑1.56 88.10↑0.34 78.12↑1.27 | 76.18↑0.54 | 76.24↑1.04 | 63.57↓1.08 61.73↑1.62 58.34↑0.60 | 69.42↓0.21 | 61.09↑0.47 |
| GPT-4-turbo | 73.03↓7.72 87.42↑1.67 77.14↓3.80 | 72.73↑0.45 | 73.94↓1.47 | 46.28↓5.78 54.42↑5.33 47.26↓1.64 | 58.32↓1.90 | 48.28↓0.74 |
| GPT-4o-mini | 76.23↓0.30 67.71↑0.31 67.92↓0.03 | 29.08↓1.00 | 34.24↓0.97 | 33.75↑0.48 45.07↑0.52 36.51↑0.52 | 27.74↓0.41 | 24.64↓0.21 |
| Claude 3 Opus | 65.60↑2.21 85.01↓4.57 72.07↓1.31 | 75.20↓2.19 | 72.09↓2.45 | 51.10↑1.03 74.59↓1.13 58.03↓0.91 | 69.95↓2.41 | 60.72↓1.29 |
| Claude 3.5 Sonnet | 80.36↑2.42 87.93↓0.62 82.33↑0.46 | 74.93↓5.30 | 77.12↓4.45 | 40.74↑5.84 65.18↓0.79 47.75↑3.51 | 58.55↑1.29 | 48.83↑2.60 |
| Claude 3.5 Haiku | 62.69↓8.39 76.01↑0.37 66.02↓5.66 | 40.35↓9.13 | 46.02↓10.1 | 37.42↓6.90 48.72↓4.65 39.91↓6.63 | 29.64↓5.71 | 27.22↓4.48 |
| Gemini 1.5 Pro | 84.64↓4.38 82.25↑4.43 79.73↑1.36 | 76.86↓0.53 | 76.71↑0.49 | 50.40↓4.13 52.33↑0.28 48.42↓2.11 | 69.09↓6.08 | 54.83↓4.45 |
| Llama-3.1-405B | 79.09↑2.72 77.96↓1.60 75.16↑0.36 | 64.25↓0.01 | 67.58↑0.01 | 38.19↓1.08 58.47↑7.38 43.39↑1.12 | 52.98↓2.44 | 43.79↑0.11 |
| Mistral Large | 68.81↑1.40 78.55↓0.18 71.07↑0.89 | 60.82↓2.62 | 63.63↓1.59 | 53.91↓3.05 56.78↓0.32 52.57↓1.94 | 46.24↓1.71 | 46.36↓2.26 |
| DeepSeek-V3 | 86.98↓1.19 73.83↑1.37 75.66↑1.07 | 64.98↑1.84 | 66.50↑1.61 | 47.64↑0.65 42.12↑0.60 42.17↑0.65 | 56.02↑0.77 | 44.09↑0.72 |
| Qwen-Max | 74.59↓5.53 79.86↓1.30 73.79↓2.78 | 60.93↓1.24 | 64.34↓1.59 | 46.49↓0.02 53.59↓0.35 46.87↓0.17 | 45.29↓1.54 | 40.87↓0.23 |
| GLM-4-Plus | 80.28↓1.03 72.81↑0.61 71.17↑0.34 | 69.34↓1.73 | 67.55↓0.43 | 43.50↓1.77 46.93↑0.43 41.64↓0.61 | 55.80↓3.17 | 43.54↓1.26 |
5 Experimental Setup
System Selection.
We select 12 LLMs as summarizers, covering the most advanced and lightweight variants777Our evaluation began on January, 2025, and all LLMs used the latest official API versions available at that time.: GPT-4o, GPT-4-turbo, GPT-4o-mini, Claude 3 Opus, Claude 3.5 Sonnet, Claude 3.5 Haiku, Gemini 1.5 Pro, Llama-3.1-405B, Mistral Large, DeepSeek-V3, Qwen-Max, GLM-4-Plus. All model sources are listed in Appendix E.
Prompt Engineering.
We benchmark LLMs for KGDS under two interaction patterns: (1) Single-turn Structured-prompt: A well-structured standard prompt Li et al. (2023) that includes input content, input definition, task description, output definition, and return format (Appendix F.3). (2) Multi-turn Self-reflection: A second-round self-reflection instruction Shinn et al. (2023) with step-by-step chain-of-thought reasoning Wei et al. (2022) following the structured prompt (Appendix F.4).
@@ 5 Experimental Setup
We use GPT-4o888Version: gpt-4o-2024-11-20 to perform fact and opinion verification as well as error detection due to its excellent consistency with human judgment Song et al. (2024a, b). All prompts are provided in Appendices F.5 and F.6.
@@ 6 Analysis
In this section, we primarily reveal the challenges LLMs face in KGDS and analyze the commonalities and differences among them. Sections §6.1, §6.2, and §6.3 respectively discuss the performance in background summary, opinion summary, and both paradigms under the structured-prompt setting. §6.4 explores the impact of LLM self-reflection.
6.1 Background Summary
@@
From Table 3, we find that most LLMs achieve moderate retrieval performance () and lightweight models (i.e., GPT-4o-mini and Claude 3.5 Haiku) perform poorly. However, from and , we observe significant polarization among LLMs, which indicates distinct retrieval strategies: some prioritize precision at the cost of recall (e.g., GPT-4-turbo), while others do the exact opposite (e.g., DeepSeek-V3). Such imbalance reveals that the current LLMs struggle with the precision-recall trade-off. Moreover, we investigate the paragraph retrieval ratio999The ratio of the number of retrieved paragraphs to the total number of paragraphs. (Figure 3) and identify that most LLMs exhibit either under- or over-retrieval, which is consistent with imbalance.
@@
As presented in Table 3, all LLMs exhibit unsatisfactory performance (). The low (average of 46.08%) reveals that LLMs often omit key facts, while the weak (average of 54.99%) indicates persistent inclusion of irrelevant facts. This dual-failure reflects the fundamental deficiencies of LLMs in meeting the requirements of coverage and focus. We also observe polarization among LLMs: some prioritize precision at the cost of recall (e.g., Claude 3.5 Sonnet), while others attempt to balance both (e.g., GPT-4o). Unlike EBS, we do not find any extreme recall-oriented models, indicating that LLMs tend to be either conservative or balanced in ABS generation.
6.2 Opinion Summary
@@
From Table 3, we find that decreases as the background summary quality declines (i.e., ) across all LLMs when switching paradigms. Furthermore, as shown in Figures 4 and 5, is highly positively correlated with and among LLMs in both independent paradigms. These findings indicate that a high-quality background summary facilitates opinion integration. Meanwhile, individual differences suggest that the model-inherent integration ability is also a key factor affecting AOS performance. For instance, Claude 3.5 Sonnet and Gemini 1.5 Pro show relatively weaker (i.e., below the line in Figure 4) and stronger (i.e., above the line in Figure 5) integration capabilities in EBS-AOS and ABS-AOS paradigms, respectively.
Nevertheless, even the most advanced models achieve only an average of 72% across both paradigms, indicating that numerous opinions are being incorrectly integrated. Moreover, lightweight models (i.e., GPT-4o-mini and Claude 3.5 Haiku) exhibit significant performance flaws, suggesting that opinion integration is sensitive to model scale.
AOS error distribution analysis.
As illustrated in Table 5, a significant ratio of errors is concentrated in IRU and IRIC across all LLMs under both paradigms, indicating that LLMs struggle to clarify implicit references effectively and correctly during opinion integration. Moreover, although the ratios of OFI, OSD, and OM errors are relatively lower, their presence still highlights inherent deficiencies such as factual inconsistency, sentiment distortion, and incorrect participant assignment in LLMs.
6.3 Summarization Paradigm
Performance stratification among LLMs.
We observe that the overall metric exhibits distinct hierarchies in independent paradigms and cross-paradigm (Figures 6, 8, and 9 in Appendix A). This suggests that the performance of LLMs improves in a stepwise manner rather than continuously as their intelligence advances in KGDS. Nevertheless, even the best-performing models achieve less than 69% average performance across both paradigms (see Figure 9), highlighting that KGDS remains a challenging task for current LLMs.
Our evaluation framework aligns well with human judgment.
To validate this, we conduct a human evaluation of overall paradigm-level quality101010We focus on paradigm-level validation because the KGDS task’s success depends on the complementarity of both sub-summaries. Furthermore, the individual sub-summary evaluations are already grounded in expert annotations and interpretable, fine-grained verification methods that are known to correlate well with human assessment Song et al. (2024a).. We randomly sample 120 summary pairs (10 per model and 5 per paradigm) from all outputs, and two human evaluators rate their overall quality on a 1-5 Likert scale Joshi et al. (2015), using the average as the final human score. For comparison, we select three LLMs (i.e., GPT-4o, Llama-3.1-405B, DeepSeek-V3) as LLM-as-a-judge baselines to directly score the summaries, with prompt settings following G-Eval Liu et al. (2023a). Table 4 presents the Pearson, Spearman, and Kendall correlation coefficients. Our framework achieves significantly higher correlation with human judgments than the baselines, demonstrating the effectiveness of our fine-grained, hierarchical approach.
| Metric | Pearson | Spearman | Kendall |
| LLM-as-a-judge | |||
| GPT-4o | .5273 | .5162 | .4193 |
| Llama-3.1-405B | .3506 | .3760 | .2881 |
| DeepSeek-V3 | .4282 | .4417 | .3479 |
| Our Framework | .6717 | .6602 | .5829 |
@@ 6.4 Self-Reflection Impact
From Table 3, we observe that the performance fluctuations (i.e., or ) after self-reflection are limited (the maximum average fluctuation across all metrics is for ). This means that self-reflection does not fundamentally influence the capacities of LLMs in KGDS, revealing two key limitations: (1) LLMs lack sufficient self-evaluation abilities for KGDS. (2) The reflection strategies of LLMs struggle to provide excellent reasoning paths for KGDS.
@@
As presented in Table 5, most LLMs exhibit a decrease in IRIC and an increase in IRU (e.g., GPT-4o with IRIC , IRU ). This suggests that self-reflection encourages LLMs to adopt more cautious strategies for clarifying implicit references — opting to leave them unclarified rather than risk incorrect clarification. However, such conservatism ultimately harms the quality of opinion summaries, as evidenced by the widespread drop in in Table 3.
7 Related Work
Dialogue Summarization. This task spans diverse scenarios, including online chats Gliwa et al. (2019), meetings Hu et al. (2023), and customer service Lin et al. (2021). Existing methods, whether based on feature modeling Chen and Yang (2021); Fang et al. (2022); Lin et al. (2022), pre-training Zou et al. (2021); Zhong et al. (2022), or LLMs Wang et al. (2023a); Zhu et al. (2025), focus heavily on single-source dialogue inputs while neglecting shared background knowledge among participants. Moreover, current evaluation systems Ramprasad et al. (2024); Tang et al. (2024b) also assume single-source inputs. In contrast, our KGDS focuses on multi-source inputs and complementary outputs.
Knowledge-Intensive Dialogue Response. This task aims to enhance the dialogue process by incorporating external knowledge sources Lewis et al. (2020); Chen et al. (2024). It differs from our KGDS in two key aspects. First, knowledge-intensive dialogue focuses on improving the conversational experience for participants Wang et al. (2023b, 2024), whereas KGDS is designed to generate clear summaries for external readers. Second, the former addresses knowledge gaps between participants, making the dialogue inherently knowledge-intensive Zhang et al. (2020). In contrast, KGDS centers on discussions among participants with shared knowledge, resulting in background-sparse conversations that rely heavily on implicit context.
Summarization Evaluation. Earlier similarity-based metrics like Rouge Lin (2004), BERTScore Zhang et al. (2019), and MoverScore Zhao et al. (2019) are often misaligned with human judgment. Recently, LLM-based evaluators Chen et al. (2023); Shen et al. (2023); Fu et al. (2024) have been used, yet they still lack interpretability. To address this, some studies Song et al. (2024a); Lee et al. (2024); Yang et al. (2024); Scirè et al. (2024) utilize atomic facts Min et al. (2023) and LLM-based claim verification Song et al. (2024b), achieving more fine-grained evaluation. Among these, FineSurE Song et al. (2024a) is most relevant to our work. However, its metrics lack consistency in evaluation granularity. In contrast, our coverage and focus metrics are both measured at the atomic fact level, ensuring consistent granularity across dimensions.